OpenStudy (anonymous):

Engine oil (η = 0.20 Pa⋅s) passes through a fine 1.90-mm-diameter tube that is 9.40cm long. What pressure difference is needed to maintain a flow rate of 6.4mL/min ?

OpenStudy (anonymous):

I did, Q = (Pi)( R^4) (delta P) / 8 (flow rate )(length) P= 8 (1.067x10^-4) (0.20) (0.094) / (Pi) (9.5x10^-4) = 6268625 Pa

OpenStudy (anonymous):

this is wrong i dont undetstand why? and what did I do wrong?

OpenStudy (anonymous):

can someone please help me, I have been stuck on this one problem for an hour, pleasee

OpenStudy (anonymous):

@Luigi0210